How much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 60% antifreeze?
Answer by stanbon(48558)

How much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 60% antifreeze?
---------------
Antifreeze DATA:
Amount = 1 gallon :Amt of active ingredient = 1 gallon
---------------
Water DATA:
Amount = x gallons ; Amt of active ingredient = 0 gallons
-----------------
Mixture DATA:
Amount = x+1 gallons ; Amt of active ingredient = 0.60(x+1) = 0.6x + 0.6
-----------------
EQUATION:
active + active = active
1 + 0 = 0.6x+0.6
0.4 = 0.6x
x = (2/3) gallon of water must be added
===================
Cheers,
Stan H.